OpenStudy (anonymous):

Equivalent Expressions 11-u-14u +13u = 11 + 17u Show work and explain how you do please. <3

OpenStudy (anonymous):

11-u-14u+13u=11+17u first, combine like terms: 11-2u=11+17u then subtract 11 from one side: -2u=17u add -2 u to 17u= 19u=0 u=0